OpenStudy (anonymous):

Mega Movies hosted a film premiere on Friday night. They charged $9 for adults and $4 for children. One hundred fifteen adults and children attended, and $940 was made in ticket sales. How many children and how many adults went to the film premiere? a)16 children; 99 adults b)23 children; 92 adults c)13 children; 102 adults d)19 children; 96 adults

OpenStudy (mathmate):

@madmoiselle_sarcasm Would you like to do this in your head or like to solve system of two equations?

OpenStudy (anonymous):

Solve the system

OpenStudy (mathmate):

Then we'll figure it out the long way: Can you first set-up/define the variables? You can use x and y, or c and a. I like the latter.

OpenStudy (anonymous):

c+a=115 $9a+$4c=$940

OpenStudy (mathmate):

So now's the time to solve the system. How far have you gone?

OpenStudy (anonymous):

9a+4(115-a)=940

OpenStudy (mathmate):

Great! You're using the substitution method. Excellent. Please continue.

OpenStudy (anonymous):

9a+460-4a=940

OpenStudy (anonymous):

9a-4a=480 5a=480 a=96

OpenStudy (anonymous):

That's right

OpenStudy (mathmate):

Formidable, madmoiselle! Could you continue to get c?

OpenStudy (anonymous):

9*96+4c=940 864+4c=940 4c=76 c=19

OpenStudy (mathmate):

Perfect, you got the correct answer! Well done!

OpenStudy (anonymous):

Thank you!

OpenStudy (mathmate):

You're welcome! À la prochaine!